three years ago, Rica was thrice as old as Dandy. Three years hence, she will be twice as old. find their present ages.
Answer by sachi(392)

three years ago, Rica was thrice as old as Dandy. Three years hence, she will be twice as old. find their present ages.
ans:-
let Dandy is x yrs now ,so 3 yrs ago he was x-3 yrs
so rica was 3(x-3)yrs then
3yrs hence dandy will be x+3 & rica will be 3(x-3)+6
as per question 3(x-3)+6=2(x+3)
or 3x-9+6=2x+6
or x=9= dandy;s age now
& rica is 3(9-3)+3=21 yrs ans
otherwise
let dandy is x yrs & rics is y yrs now
three years ago y-3=3(x-3)
or 3x-y=6----------1
Three years hence y+3=2(x+3)
or 2x-y=-3 --------2
subtracting (2) from(1)
x=9
from (1) y=21 ans